F17 UTA-701 Measurement of Volume. Complete the following calculations. a. 50 ml, beaker 2. #1 #2 #3 % e2 sees 56.6y a, 1 . Mass of beaker with -30mL ofwater a.2. Mass of empty beaker a.3. Mass of water a.4. True volume in beaker (calculated from mass and density) a.5. Observed Volume (as read from beaker)9 m a.6. Percent (%) error a7. Average % error Because each measurement was made with a slightly different amount of water, we will normalize the three measurements to a true expression: volume of 30.00 mL using the following Where Vso is 30.00 mL, V, is true volume (2.a 4) as determined by massi/density, Vas is the observed volume (a. S above) and V, is the normalized observed volume. ?.8. Normalized Observed Volume a.9. Average V, (avg of 2a.8) a.10. Average absolute deviation (calculated using Avg. V and three individual V) 24
Answer #1

Mass of water = 56.64-28.53 , 56.64-28.53, 56.63-28.53 (2a.1-2a.2)

                       = 28.11, 28,11 , 28.1 grams

Volume = mass/density = 28.11, 28.11, 28.1 ml (density = 1gm/ml)

V30 = Vobs = 30ml

Normalized volume = V30xVobs/V1 = 32.01 ml,32.01ml, 32.02 ml

Average Vn = (32.01+32.01+32.02)/3 = 32.013 ml

Deviation = Normalized Vn - Average Vn = -0.003,-0.003,0.007 ml

Average deviation = (-0.003-0.003+0.007)/3 = 3x10-4 which is approximately zero.